" Given: " f(x) = 3x - 5 " and " g(x) = -5x + 2
1. " Find " f[g(-1)]
2. " Find " g[f(4)]

Answer by stanbon(75887)

" Given: " f(x) = 3x - 5 " and " g(x) = -5x + 2
1. " Find " f[g(-1] = f(-5*-1+2) = f(7) = 3*7-5 = 16
------------------------------------
2. " Find " g[f(4)] = g(3*4-5) = g(7) = -5*7+2 = -33
===========================================================
Cheers,
Stan H.
==========